I'd greatly appreciate it if some one could help me. Thanks!! Richard's average driving speed is 7 miles per hour faster than Mia's. In the same length of time it takes Richard to drive 318 miles, Mia only drives 276 miles. What's Richard's average driving speed.
Answer by edjones(8007)

d=st d=dist, s=speed, t=time
Let Mia's speed be s.
t=d/s
318/(s+7)=276/s
318s=276(s+7)
318s=276s+1932
42s=1932
s=46 mph
46+7=53 mph Richard's speed
.
Check:
318/53=6 hrs
276/46=6 hrs
.
Ed